OpenStudy (anonymous):

Solve this problem. n - 16 = 2n + 12 + 3n

OpenStudy (anonymous):

so we are solving for n. just like ther other one, combine like terms on that right side.

OpenStudy (anonymous):

so it would be n - 16 = what

OpenStudy (anonymous):

n-16=2n+3n right

OpenStudy (anonymous):

almost exactly. n - 16 = 2n + 3n +12

OpenStudy (anonymous):

so to get rid of that n, we want to get all the n's on one side. if you want to minus the n from the left or minus the 5n from the right, you can do either. type out what you get

OpenStudy (anonymous):

i dont get you

OpenStudy (anonymous):

alright. let me say it another way

OpenStudy (anonymous):

now you just solved the combining like terms part to be n-16 = 2n+3n+12, which is n-16 = 5n+12. so now we have to get both the different terms on different sides: the n's on one side and the regular numbers on the other side.

OpenStudy (anonymous):

to do this we have to subtract from both sides of the equation to get on one side, we will start with the n's....... we will take the n from the left and move it to the right so it will be written as:

OpenStudy (anonymous):



OpenStudy (anonymous):

so what do we do to move the 12 over to the left hand side?

OpenStudy (anonymous):



OpenStudy (anonymous):

and then just solve for n. -28 / 4 = n